QUADRATIC ORNOT QUADRATIC? The Length of a swimming pool is 8m longer than its width and the areais 105m2. don't solve it write the problem in a mathematical sentence thanks :)
Answer by ankor@dixie-net.com(22740)

The Length of a swimming pool is 8m longer than its width and the area is 105m2. don't solve it write the problem in a mathematical sentence
:
Let x = width
then
(x+8) = the length
the area, length * width
x(x+8) = 105
x^2 + 8x - 105 = 0, a quadratic